Question 12 of 12 What is the solubility of Mg(OH)2 at a pH of 11.70? (Ksp Mg(OH)2 is 1.6 * 10-13) | 4 5 6 c X 100
Answer #1

Мо бо ), a Mat2 +2° s 25 wheh, b = 1-75 , A4 25°С, Бек = 1ч - Ен = ч- .70 = 220 А4, he = - log [ои:) е) [) = lo-23 = 0. oos M